there are two electric bulbs (1) marked 60 w,220 v (2) marked 100 w,220 v. which one of them has higher resistance?

Grade:10

1 Answers

Vikas TU
7 years ago
R = v^2/P
R1 = 220*220//60 = 806.66 ohm.
R2 = 220*220/100 = 484 ohm.
Now U can compare easily.
Bulb 2 has more resistance.